OpenStudy (anonymous):

sec^2x/sec^2x-1=csc^2x

OpenStudy (anonymous):

Do you need to verify the identity?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, first you need to change the form of the left side using fundamental trig identities. The important trig identities you need are:\[secx=\frac{ 1 }{ cosx}\]\[\sec^2x-1=\tan^2x\]

OpenStudy (anonymous):

\[\frac{ \sec^2x }{ \sec^2x-1 }=\csc^2x\] Since secx=1/cosx, sec^2x=1/cos^2x. Let's substitute this as well as the second identity.\[\frac{ 1 }{ \cos^2x }*\frac{ 1 }{ \tan^2x }\]

OpenStudy (anonymous):

All we've done so far is change the form of the left side of the equation. Another important identity is \[tanx=\frac{ sinx }{ cosx }\] We can use this identity to change the form of 1/tan^2x. Since tan^2x is on the bottom, we have to take the reciprocal of the identity. This is what we have so far. \[\frac{ 1 }{ \cos^2x }*\frac{ \cos^2x }{ \sin^2x }\]

OpenStudy (anonymous):

The cos^2x cancel out so we're left with \[\frac{ 1 }{ \sin^2x }\] which is equal to\[\csc^2x\]. We have successfully verified the identity. When you do problems like these, make sure you use your trig identities to change the form of each side. Does that make sense?

OpenStudy (anonymous):

yes that makes sense thank you :)

OpenStudy (anonymous):

Of course! :)